# deepSURF: Detecting Memory Safety Vulnerabilities in Rust Through Fuzzing LLM-Augmented Harnesses

### v1

@@ Appendix B Harnesses Selector Policies
TABLE IV: Bug detection capability of deepSURF without LLM-based harness augmentation. Each row corresponds to a unique bug. ✓ indicates detection via fuzzing, ✗ indicates failure to detect. * signifies a new bug discovered by deepSURF; otherwise, the assigned RustSec ID is provided.
Appendix C Results of the Ablation Study
To assess how deepSURF’s core components contribute to its effectiveness in detecting memory safety bugs and fuzzing unsafe code reachable through a library’s API, we conducted an ablation study targeting three key components: (1) static analysis for static harness generation, (2) LLM-based augmentation, and (3) the harness selection policies. We compared the full version of deepSURF against four alternative configurations, each disabling or modifying one of these components, as summarized in Table V.
We evaluated these configurations on three crates from the ERASAN dataset that contain memory safety bugs requiring harnesses with diverse characteristics to be exposed. Since each configuration yields a different number of harnesses, and to ensure a fair comparison, we allocated the same total fuzzing time per configuration: 672 CPU-hours. Thus, the fuzzing time per harness is computed as:
[…]
This approach gives more fuzzing time per harness to configurations that yield fewer harnesses. We fuzz each target using two AFL++ threads in a master-slave setup: one with ASan and the other with CmpLog.
To measure unsafe code coverage, we count code lines within unsafe blocks or unsafe functions that are reachable through safe APIs of the evaluated crates. We exclude empty lines, comments, and feature-gated code not compiled under the default feature set.
Configuration Components
Static Harness
Generation
LLM-based
Augmentation
Harness Selection Policy
deepSURF
Skip URAPIs without
custom functionality
deepSURF-static
deepSURF-llm
deepSURF-skip-all Skip any URAPI
deepSURF-no-skip
Do not skip
C.1 Static vs. LLM-based Analysis
We distinguish the deepSURF-static and deepSURF-llm configurations to evaluate the individual contributions of static analysis and LLM integration to deepSURF’s ability to fuzz unsafe code and uncover memory corruption bugs.
Crate deepSURF deepSURF-llm deepSURF-static
Detected
Bugs
Unsafe
Coverage
URAPI Coverage
(#Harnesses)
Detected
Bugs
Unsafe
Coverage
URAPI Coverage
(#Harnesses)
Detected
Bugs
Unsafe
Coverage
URAPI Coverage
(#Harnesses)
toodee 4 83.8% 90% (35) 0 47.2% 35% (9) 3 83.1% 81.7% (159)
simple-slab 2 100% 100% (3) 2 100% 100% (12) 0 50% 100% (14)
smallvec-1.6.0 1 86.8% 84.7% (56) 0 85.9% 66.7% (44) 0 0% 0% (0)
TOTAL 7 86.2% 87.9% (94) 2 72.2% 56.4% (65) 3 32.7% 40.7% (173)
(a) Impact of static analysis and LLM integration.
Crate deepSURF deepSURF-skip-all deepSURF-no-skip
Detected
Bugs
Unsafe
Coverage
URAPI Coverage
(#Harnesses)
Detected
Bugs
Unsafe
Coverage
URAPI Coverage
(#Harnesses)
Detected
Bugs
Unsafe
Coverage
URAPI Coverage
(#Harnesses)
toodee 4 83.8% 90% (35) 2 92.9% 91.7% (31) 3 95.8% 95% (49)
simple-slab 2 100% 100% (3) 2 100% 100% (3) 2 100% 100% (9)
smallvec-1.6.0 1 86.8% 84.7% (56) 0 81.5% 81.9% (22) 1 89.4% 84.7% (49)
TOTAL 7 86.2% 87.9% (94) 4 84.9% 87.1% (56) 6 89.1% 90% (107)
TABLE VI: Ablation study results across different deepSURF’s configurations.
Based on our results, deepSURF achieves the highest bug-finding capability (seven bugs) and the highest unsafe code coverage (86.72%) compared to the other two configurations, demonstrating that the combination of static analysis and LLM-based augmentation outperforms approaches that rely on only one of these components.
We observed that deepSURF-llm struggled to identify URAPIs, particularly in the case of toodee, where it returned APIs that could not reach unsafe code and failed to generate harnesses for 65% of the crate’s URAPIs. Although it achieved high unsafe code coverage in smallvec-1.6.0, it did not trigger the bug, which required custom trait implementations—a feature supported by deepSURF through its static analysis. In contrast, for simple-slab, deepSURF-llm performed comparably to deepSURF, discovering both memory safety bugs. Upon inspection, we found that simple-slab has a relatively simple API (e.g., lacking complex trait relationships), and its bugs depend primarily on sequences of API calls, a feature supported in both deepSURF and deepSURF-llm.
On the other hand, deepSURF-static detected only three bugs in toodee and achieved much lower unsafe code coverage than the other configurations. Its success in toodee is attributed to static control flow analysis and support for custom trait implementations. However, it failed to find any bugs in simple-slab due to the lack of sequence support. Also, it could not find the bug in smallvec-1.6.0, where all URAPIs require generic arguments implementing the unsafe Array trait, feature supported only by the LLM-based augmentation. Consequently, it was unable to generate valid harnesses for smallvec-1.6.0.
C.2 Comparing Harness Policies
In Section 4.4, we described the operation of the Harness Selector and Harness Augmenter, along with their corresponding harness selection policies. These policies influence both the composition and size of the final Fuzzing Corpus and, consequently, affect deepSURF’s bug-finding capability. In this section, we evaluate alternative harness selection policies for both components.
The results of this comparison are presented in Table VI(b). Among all harness selection policies, deepSURF’s default policy achieves the highest bug-finding capability. The second-best configuration is deepSURF-no-skip, which detects six bugs. The only missed bug is one that deepSURF identifies using a statically generated harness containing custom trait implementations. As discussed earlier, deepSURF-no-skip does not retrieve such harnesses from the static generation stage.

### v2

@@ Appendix A Harnesses Selector Policies
Appendix B False Positives Analysis
As noted in Section 2, fuzzing Rust harnesses can yield crashes that are not genuine bugs (false positives). deepSURF targets automatic detection of memory safety bugs in Rust libraries while keeping manual triage effort low. An output dominated by false positives would undermine the tool’s utility. To mitigate this, deepSURF configures afl.rs to ignore panics that do not lead to memory corruption (see Section 4.5). In this section, we empirically evaluate the effectiveness of this design choice.
We introduce deepSURF-no-filter, a configuration of deepSURF that keeps panic hooks enabled (see Section 4.5). We fuzz both deepSURF and deepSURF-no-filter using the settings and crates of our ablation study (see Section 6). deepSURF produces 630 total crashes; deepSURF-no-filter produces 5,764. Leveraging ASan’s instrumentation which forces a crash when memory violation occurs during fuzzing, both variants auto-classify memory corruption crashes—607 for deepSURF and 292 for deepSURF-no-filter—all true positives. These deduplicate into seven and six distinct memory safety bugs, respectively. After removing the ASan-confirmed cases, 23 crashes (deepSURF) and 5,472 crashes (deepSURF-no-filter) remain for manual triage. For deepSURF, all 23 are false positives and can be dismissed within minutes of human effort. For deepSURF-no-filter, 5,118 of the remaining crashes are false positives, requiring at least a day of manual work to classify.
Moreover, deepSURF-no-filter triggers 354 crashes attributable to 12 distinct arithmetic bugs (e.g., addition overflows) in the fuzzed crates. Because afl.rs enables arithmetic overflow checks by default, these cases appear as panics during fuzzing. Notably, four of these arithmetic bugs occur inside unsafe blocks and could potentially lead to memory corruption when building in release mode, where overflow checks are typically disabled. By design, deepSURF misses these 12 bugs (therefore, they can be considered as false negatives), due to the fact that in its current version panics from arithmetic overflows are ignored. However, it could be configured to detect the ones that manifest as memory bugs by compiling the harnesses with the build setting overflow-checks=off.
Appendix C Meta-Review
The following meta-review was prepared by the program committee for the 2026 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
C.1 Summary
This paper presents deepSURF, a system that automatically generates fuzzing harnesses for Rust libraries, using static analysis augmented with LLMs, for the purpose of finding memory-safety bugs. In an evaluation over 63 real-world crates, deepSURF covers 87.3% of URAPIs and rediscovers 30 of 56 known bugs, plus 12 new memory-safety bugs, outperforming existing Rust fuzzers.
C.2 Scientific Contributions
  • Provides a Valuable Step Forward in an Established Field
C.3 Reasons for Acceptance
  1. 1.
    The paper provides a valuable step forward in an established field. Compared to prior Rust fuzzing tools, deepSURF fuzzes more of the target library (and finds more memory-safety bugs) by generating closures for APIs that require them, and invoking related APIs to create different states when invoking target URAPIs; these advances are made possible by a creative combination of static analysis and LLMs. The paper’s empirical evaluation is well done.
C.4 Noteworthy Concerns
  1. 1.
    The paper’s evaluation presents a strong quantitative result, but could do more to readily show the effectiveness of deepSURF by qualitatively analyzing the generated harnesses and systematically comparing them to actual usages of the targeted APIs, to assess realism.
Appendix D Response to the Meta-Review
A preliminary qualitative analysis indicates that deepSURF’s generated harnesses reflect realistic usage patterns and adhere to the documented APIs and semantics of the evaluated crates. For example, for crates implementing data structures such as toodee, smallvec, slice_deque, simple-slab, and stack_dst, deepSURF produces harnesses that naturally follow common user workflows: constructing the data structure, populating it with insertion operations, and subsequently performing stateful transformations such as removals, cloning, swapping, and filtering. However, conducting a systematic qualitative assessment of the generated harnesses (whether through manual review or automated analysis) would require substantial additional effort. We will explore this as part of future work.